# Rotating axionic AdS4 black hole dressed with a scalar field

### v1

@@ Acknowledgements.
[…]
Appendix A Equations of motion
In this section, we present the explicit expression for the equations of motion given in (3) together with the line element (6) for $\xi=\frac{1}{6}$ and $N(r)=1$, which are the following:
@@ Appendix A Equations of motion
-- $\frac{1}{3}f\phi\phi^{\prime\prime}-\frac{2}{3r}\phi\phi^{\prime}f+U(\phi)+\Lambda=0,$
$\mathcal{E}^{t}_{\phantom{a}\theta}$ $=$ $2(N^{\theta})^{\prime}r-\frac{1}{3}r(N^{\theta})^{\prime}\phi^{2}+\frac{1}{2}r^{2}(N^{\theta})^{\prime\prime}-\frac{1}{12}r^{2}(N^{\theta})^{\prime\prime}\phi^{2}-\frac{1}{6}\phi\phi^{\prime}r^{2}((N^{\theta}))^{\prime}=0,$
@@
-- $\frac{1}{6}N^{\theta}\phi^{\prime}\phi f^{\prime}-\frac{N^{\theta}f}{r^{2}}+\frac{1}{6r^{2}}N^{\theta}f\phi^{2}=0,$
@@
-- $\frac{1}{6r^{2}}f\phi^{2}+\frac{f}{r^{2}}+U(\phi)+\Lambda=0,$
@@
-- $\frac{1}{3r}\phi\phi^{\prime}f-\frac{1}{3}\phi\phi^{\prime}f^{\prime}+\frac{f^{\prime}}{r}-\frac{1}{6r}f^{\prime}\phi^{2}+U(\phi)+\Lambda=0,$

### v2

@@ Acknowledgements.
The authors would like to thank the anonymous referees for carefully reading our manuscript and giving valuable suggestions that led to an improved version of this work. The authors gratefully acknowledge the insightful comments of Professor Julio Oliva. M.B. is supported by proyecto interno UCM-IN-25202. JAHM and DFHB are also grateful to SNII from CONAHCYT.
Conflict of Interest
The authors declare no conflict of interest.
Data Availability Statement
Data sharing is not applicable to this article, as no data sets were generated or analyzed during the current study.
Apêndice A Equations of motion
In this section, we present the explicit expression for the equations of motion given in (3) together with the line element (6) for $\xi=\frac{1}{6}$ and $N(r)=1$. The non-vanishing components read:
@@
- $\frac{1}{3}f\phi\phi^{\prime\prime}-\frac{2}{3r}\phi\phi^{\prime}f+U(\phi)+\Lambda,$
$\mathcal{E}^{t}_{\phantom{a}\theta}$ $=$ $2(N^{\theta})^{\prime}r-\frac{1}{3}r(N^{\theta})^{\prime}\phi^{2}+\frac{1}{2}r^{2}(N^{\theta})^{\prime\prime}-\frac{1}{12}r^{2}(N^{\theta})^{\prime\prime}\phi^{2}-\frac{1}{6}\phi\phi^{\prime}r^{2}((N^{\theta}))^{\prime},$
@@
- $\frac{1}{6}N^{\theta}\phi^{\prime}\phi f^{\prime}-\frac{N^{\theta}f}{r^{2}}+\frac{1}{6r^{2}}N^{\theta}f\phi^{2},$
@@
- $\frac{1}{6r^{2}}f\phi^{2}+\frac{f}{r^{2}}+U(\phi)+\Lambda,$
@@
- $\frac{1}{3r}\phi\phi^{\prime}f-\frac{1}{3}\phi\phi^{\prime}f^{\prime}+\frac{f^{\prime}}{r}-\frac{1}{6r}f^{\prime}\phi^{2}+U(\phi)+\Lambda,$